# Towards Locally Deployable Fine-Tuned Causal Large Language Models for Mode Choice Behaviour

### v1

@@ 3.1 Instant-Level LLMs Mode Choice Predictive Performance Evaluation
Figure 6 shows, for each model and dataset, the percentage change in weighted F1 when moving from zero-shot to few-shot random and from few-shot random to few-shot targeted. These relative gains expose each model’s capacity to learn from uncurated versus carefully chosen in-context examples.
@@
In the Swissmetro survey, most models improve by 4–15% when given random examples and by an additional 2–27% under targeted examples. Notably, Gemma 3 4B loses performance (–2.3%) with random prompts, a clear case of negative learning likely caused by unrepresentative support sets demonstrated by the strong rebound when supplied with curated examples. By contrast, the Stealth 1.2 7B and Stealth 1.3 7B models display substantial random-shot gains (10.4%, 1.1%) and the largest targeted-shot uplifts (17.2%, 27.3%), highlighting their surprising adaptability despite modest parameter counts and baseline performance. On the BrightWater dataset, which is characterized by high behavioural heterogeneity, Mistral 7B leads all models with a random-shot gain of 39.0%, while Stealth 1.2 7B peaks under targeted examples at 36.8%. Only the distilled variants of DeepSeek R1 show slight declines (–5.9%, –0.4%) with random prompts, suggesting that their fixed reasoning priors may sometimes misalign with diverse traveller profiles. However, once examples are curated, even these distillates recover and show substantial gains of 22.1–25.9%. In the London PMC context, where zero-shot baselines are already strong, Gemma 3 12B exhibits the most dramatic learning curve, with a 56.0% improvement under random examples, but also the largest volatility, dropping -3.4% when switching to targeted. Smaller LLaMA variants, such as Llama 3.2 1B, show negative learning under random (–10.6%) but rebound (+10.6%) with targeted examples, indicating that lightweight architectures are more sensitive to example quality but can still benefit from well-aligned support sets.
@@
When averaged across all surveys, the stealth models stand out with the largest sequential gains (18.7, 23%) from zero to random, and an additional 19% to 20.5% from random to targeted, combining strong learning potential with reduced negative learning. In contrast, the distilled models post modest average improvements (2.4–5.4% from zero to random, 6.6–7.4% from random to targeted), confirming their limited plasticity once reasoning patterns are fixed during distillation.
@@
These findings support three key insights. First, negative learning can occur when support examples conflict with pretrained expectations or when selected samples are largely misaligned with the test case, particularly in smaller models. Second, instruction-tuned, mid-sized models like those from the Stealth family offer a strong adaptability and learning potential with substantial gains once presented with in-context examples. Third, large, dense transformers such as Gemma 3 12B can leverage random support sets effectively, and learn the underlying relationship between the commuters’ profile and their mode choice, which can be further improved with curated examples. In practice, deploying a few random-shot learning style without quality control should be avoided, especially in complex or noisy environments. Where curated examples are available, their impact can be substantial. Among the tested models, Stealth variants emerge as robust and versatile learners, while distilled models remain a reliable choice for zero-shot deployment when examples cannot be supplied.
@@
Building on the earlier findings regarding model architecture, dataset characteristics, and shot styles, we now turn to a finer grained examination of two additional dimensions influencing predictive performance: the prompt style (Direct vs. Chain-of-Thought/ReAct) and the model temperature (0.5 vs. 1). These two axes offer insight into how LLMs respond to adjustments in prompt structure and sampling stochasticity, independent of parameter count or dataset complexity. These elements, while often overlooked, introduce subtle but meaningful shifts in model behaviour across the three travel survey datasets. Figure 7 illustrates, for each model across the survey dataset, the percentage point change in weighted F1 score when switching from Chain-of-Thought/ReAct to Direct prompting and from temperature 0.5 to temperature 1, across all three shot styles: zero-shot, few-shot random, and few-shot targeted.
@@
Across the board, differences arising from prompt style were larger and more systematic than those from temperature changes. Switching from a Chain-of-Thought/ReAct prompt to a Direct prompt generally improved performance once illustrative examples were available. This effect was most pronounced in the Fewshot_Targetted learning style, where Direct prompting consistently sharpened the model’s focus. For instance, with temperature held at 0.5, using a Direct prompt in the Fewshot_Targetted setting resulted in positive mean F1 score changes across all three datasets (Brightwater: +1.58%, London PMC: +4.63%, Swissmetro: +0.52%) and a positive impact on the weighted F1 score in 9 out of 11 models. However, results in the Zeroshot were noisier and more model dependent; while most models (7) still benefited from concise Direct prompts, others (4) performed better with the reasoning structure provided by CoT/ReAct. These patterns confirm that prompt format is not a one-size-fits-all hyperparameter but must be aligned with the model’s architecture and the specific task context.
@@
Temperature adjustments (from 0.5 to 1.0) produced more subtle effects. Contrary to improving performance, increasing the temperature to 1.0 often led to negligible changes or slight performance declines, especially under direct prompting style, indicating a general preference for the more deterministic decoding associated with a temperature of 0.5. For example, when using Direct prompting, the number of models that lost performance at $T=1.0$ was greater than the number that gained performance across all datasets and shot types. A critical finding is the non-trivial interaction between temperature and prompt style. The effect of changing temperature often diverged in both magnitude and direction depending on the prompt. For the Qwen 2.5 7B model, increasing temperature under a Direct prompt degraded performance by 3.5-3.5- 3.5% for few-shot targeted, whereas under a CoT/ReAct prompt, it improved performance by + 5.7%.
@@
Architectural and training differences help explain these heterogeneous responses. Dense instruction-tuned models (e.g., the Gemma, Llama, and Mistral families) typically capitalized on Direct prompts, especially in the Fewshot_Targetted setting, reflecting their fine-tuning on concise task specifications. Larger models, like Gemma 3 12B model, although showed consistent improvement with Direct prompting at $T=0.5$, it was overall less sensitive to varying prompting styles or temperatures. On the other hand, smaller models showed extreme sensitivity, with the Llama 3.2 1B model showing a massive +30.37% F1 score increase on the London PMC dataset under these conditions. In contrast, certain distilled reasoning models, such as DeepSeek R1 Distill Qwen 7B, proved more fragile to prompt compression, often losing accuracy when the CoT/ReAct reasoning chain was removed. These models were also more sensitive to increases in temperature, which often introduced harmful variability. Under CoT/ReAct prompting on the Swissmetro Fewshot_Targetted dataset, the DeepSeek R1 Distill Qwen 7B and DeepSeek R1 Distill Llama 8B models saw significant performance drops of 8.83-8.83- 8.83% and 5.01-5.01- 5.01%, respectively, when temperature was raised to 1.0.
@@ 3.2 Distribution-Level LLMs Mode Choice Predictive Performance Evaluation
Across all datasets, DistMAE and JSD exhibit a strong, roughly linear correlation because both capture how closely the aggregate selection frequencies match the empirical mode shares. Cross-entropy is correlated as well, but it is far more punitive when a present class receives zero or near-zero aggregate share in a run, which results in the large outliers.
In the Swissmetro dataset (Figure 8), models such as DeepSeek R1 Distill Llama 8B, Gemma 3 12B, and DeepSeek R1 Distill Qwen 7B achieve both high instant-level accuracy and low distribution-level error. Their selections keep minority modes represented, so no class collapses to a zero count and cross-entropy remains low. Weaker models, such as Mistral 7B and Stealth 1.3 7B, show volatile cross-entropy with visible outliers. These runs concentrate too heavily on the dominant mode and leave one or more valid classes with zero selections. The effect is most evident under zero-shot, where the absence of exemplars encourages collapse to the majority class.
Brightwater shows a different profile. Instant-level accuracy is lower, while the distribution picture is mixed: DistMAE sits in a tight band ($\approx$0.02-0.12), whereas JSD and cross-entropy vary widely, with cross-entropy spanning roughly 1.6–11 (Figure 9). Two structural features explain this. First, the choice set has nine dynamic alternatives with several low-share modes. With many thin tails, absolute deviations accumulate slowly: a model can mis-allocate small amounts across rare modes and still keep the sum of absolute errors modest, hence the low DistMAE. JSD, by contrast, responds to relative errors: a shift from 0.01 to 0.03 is small in absolute terms but large in relative terms, so JSD rises even when DistMAE stays low. Cross-entropy is stricter still: if a run assigns zero selections to a present class, the loss spikes. Second, the respondent population is comparatively homogeneous. Although the model never observes the aggregate distribution, many instances share similar covariates, so models tend to produce similar selections across cases. When aggregated, the resulting shares sit close to the empirical shares (low DistMAE) but per-instance discrimination remains modest (lower F1). Rare atypical cases then trigger confident mistakes that lift cross-entropy. Random few-shot amplifies this when exemplars are not representative; targeted few-shot counters it by anchoring selections to contextually similar cases, which lowers JSD and reins in the extreme cross-entropy values, even if the gain in F1 is smaller.
London PMC achieves the highest instant-level accuracy, with moderate distribution errors and few cross-entropy outliers (Figure 10). Three features contribute. First, the option set contains fewer very low-share classes than Brightwater, so tail-risk omissions are rarer and cross-entropy is steadier. Second, covariates separate the dominant mode clearly—time–cost trade-offs and access variables support a firmer decision boundary, which lifts per-instance accuracy and reduces the need to spread selections thinly. Third, the aggregate shares are stable enough that few-shot exemplars provide useful priors; random few-shot already attains high F1, while targeted few-shot further improves calibration by lifting minority-mode counts and lowering JSD with only small changes in F1. Even so, distribution metrics do not collapse to a floor. DistMAE and JSD remain mid-range because high-accuracy models concentrate selections on the predicted winner and trim secondary modes. Qwen 2.5 7B and Mistral 7B sit near the top in F1 while keeping DistMAE and JSD low, confident but not excessively concentrated. Stealth 1.2 7B provides a complementary profile: slightly lower F1 but the best DistMAE and among the lowest cross-entropy, indicating aggregate shares close to the ground truth. Gemma 3 12B reaches very high F1 under random shots but often places too much weight on the chosen alternative, which in some cases lifts DistMAE and JSD relative to the best calibrated models. Zero-shot remains weakest: accuracy drops and JSD rises as selections revert to broad priors that under-represent minority modes.
Discrepancies between instant- and distribution-level metrics follow a simple mechanism. The same covariate signals that drive correct top‑1 choices also guide selection frequencies, so some alignment is expected. Gaps arise in two directions. Decisive models concentrate selections; F1 climbs, but minority modes are under-selected, which raises JSD and, in the extreme, cross-entropy. Conservative models disperse selections; F1 falls, yet aggregation produces shares close to the empirical ones, lowering DistMAE. In short: sharper selection profiles favour F1; smoother profiles favour calibration. Prompting determines where a model lands on this spectrum. Targeted few-shot steers selections toward locally plausible alternatives for each instance; F1 improves and minority-mode counts rise, which usually lowers JSD and cross-entropy. Random few-shot improves F1 over zero-shot but can still misalign selections with the test context, leaving JSD elevated and occasionally producing cross-entropy spikes. Zero-shot yields both lower F1 and weaker calibration.
Model size and architecture matter in ways that reflect dataset structure. Larger models help on Swissmetro, where strong signals and class imbalance reward decisive outputs. Returns are smaller on Brightwater, where nine alternatives and thin tails punish omissions and reward calibrated dispersion. On London PMC, architectural design and training data (e.g., Qwen, Mistral) provide clear decision boundaries with stable calibration, so design can outweigh parameter count.
3.3 Pairwise Distribution-Level Analysis
@@
Brightwater shows a distinct pattern. Models generally achieve lower instant-level accuracy, yet the distribution-level picture is mixed: DistMAE sits in a tight band (about 0.02–0.12), while JSD and cross-entropy vary widely, with cross-entropy spanning roughly 1.6–11 (Figure 9). Two structural features explain this. First, the choice set contains nine dynamic alternatives, several with small shares. With many thin tails, absolute deviations accumulate slowly: a model can mis-allocate small amounts across rare modes and still keep the sum of absolute errors modest, hence the low DistMAE. JSD responds to relative error: a shift from 0.01 to 0.03 is small in absolute terms but large in relative terms, so JSD rises even when DistMAE stays low. Cross-entropy is stricter still: if a run assigns zero selections to a present class, the loss spikes. Second, the respondent population is comparatively homogeneous. Although the model never observes the aggregate distribution, many instances share similar covariates, so models tend to produce similar selections across cases. When aggregated, the resulting shares sit close to the empirical shares (low DistMAE) but per-instance discrimination remains modest (lower F1). Rare atypical cases then trigger confident mistakes that lift cross-entropy. Random few-shot amplifies this when exemplars are not representative; targeted few-shot counters it by anchoring selections to contextually similar cases, which lowers JSD and reins in extreme cross-entropy values, even if the gain in F1 is smaller.
@@
London PMC achieves the highest instant-level accuracy, with moderate distribution errors and few cross-entropy outliers (Figure 10). Three features contribute. First, the option set contains fewer very low-share classes than Brightwater, so tail-risk omissions are rarer and cross-entropy is steadier. Second, covariates separate the dominant mode clearly, time–cost trade-offs and access variables support a firmer decision boundary, which lifts per-instance accuracy and reduces the need to spread selections thinly. Third, the aggregate shares are stable enough that few-shot exemplars provide useful priors; random few-shot already attains high F1, while targeted few-shot further improves calibration by lifting minority-mode counts and lowering JSD with only small changes in F1. Despite these advantages, the distribution metrics do not collapse to a floor: DistMAE and JSD remain mid-range because high-accuracy models still concentrate selections on the predicted winner and trim secondary modes. Qwen 2.5 7B and Mistral 7B balance this well, sitting near the top in F1 while keeping DistMAE and JSD low. Stealth 1.2 7B offers a complementary profile: slightly lower F1 but the best DistMAE and among the lowest cross-entropy, indicating aggregate shares close to the ground truth. Gemma 3 12B reaches very high F1 under fewshot–random but often places too much weight on the chosen alternative, which in some cases lifts DistMAE and JSD relative to the best calibrated models. Zero-shot remains weakest: accuracy drops and JSD rises as selections revert to broad priors that under-represent minority modes.
Discrepancies between instant- and distribution-level metrics are therefore expected. Although the weighted F1 averages class-wise F1 using observed supports, it remains a top‑1 confusion‑matrix metric and does not enforce agreement between predicted aggregate shares and the empirical ones. A model can score highly on weighted F1 by excelling on majority classes while under‑selecting minority modes. This elevates JSD and, when a present class receives zero selections, produces large cross‑entropy. Conversely, a model that disperses selections more evenly may forgo some top‑1 hits (lower weighted F1) yet produce aggregate shares closer to the truth, which lowers DistMAE. In short: sharper selection profiles favour weighted F1; smoother profiles favour calibration.
@@
The prompting strategy controls this balance at the instance level. Targeted few‑shot steers selections toward locally plausible alternatives; F1 improves and minority‑mode counts rise, which usually lowers JSD and cross‑entropy. Random few‑shot improves F1 over zero‑shot but can still misalign selections with the test context, leaving JSD elevated and occasionally producing cross‑entropy spikes. Zero‑shot yields both lower F1 and weaker calibration.
As for the architecture families, all candidates are decoder‑only transformer chat models. Observed differences arise from pre‑training mix, instruction tuning, and alignment choices rather than from fundamentally different network topologies. On London PMC, Gemma 3 12B achieves the highest instant‑level accuracy and under fewshot–random also the lowest DistMAE and JSD, showing that strong accuracy can coincide with good calibration. Under fewshot–targeted, the Stealth 1.x (7B) models yield the lowest DistMAE and cross‑entropy with only a small drop in F1, which suits share forecasting objectives. Qwen 2.5–7B and Mistral 7B provide the most balanced profiles across shot types: high F1 with consistently low and stable distribution errors, making them robust defaults when both objectives matter.
@@
Model size effects. Figure 11 and Table 5 jointly summarise how parameter count relates to performance. The figure shows points for individual runs and per‑dataset mean trends at each parameter size. The table encodes the dominant direction of change with arrows: up ($\uparrow$), down ($\downarrow$), approximately flat ($\rightarrow$); slanted arrows ($\nearrow$,\,$\searrow$ , ↘) indicate weaker tendencies.
Second, interaction with prompting. Zeroshot exposes raw inductive bias: when the class structure is skewed (Brightwater), larger models amplify that bias unless guided. Random few‑shot often lifts F1 but does not guarantee coverage of rare modes, so JSD and cross entropy can remain elevated. Targeted few‑shot is the setting where scale reliably converts to better calibration: minority‑mode counts rise, DistMAE and JSD tilt downward, and cross‑entropy spikes disappear.
Third, training recipe versus size. Differences among families matter as much as parameter count on London PMC and, to a lesser extent, on Swissmetro under few‑shot. Qwen 2.5–7B and Mistral 7B achieve high F1 with steady calibration across shots, while Gemma 3 12B reaches very high F1 but can trim mass in some runs from secondary modes, which raises DistMAE and JSD relative to the best‑calibrated runs.
Reasoning‑distilled models. DeepSeek R1 Distill models (Llama 8B, Qwen 7B) often have an edge at the same parameter scale on Swissmetro: they separate alternatives cleanly while maintaining coverage of minority modes, yielding high F1 with low DistMAE and JSD and few cross‑entropy outliers. On Brightwater the same decisiveness can overshoot in zeroshot or random few‑shot; targeted prompting restores balance and the models then sit among the best calibrated for their size. On London PMC they remain competitive but not uniformly superior: Qwen 2.5 7B and Mistral 7B deliver comparable accuracy with stable divergence, suggesting instruction tuning and data mixture shape performance at least as strongly as parameter count.
@@
BrigthWater  Fewshot–Random $\rightarrow$/$\nearrow$ / ↗ $\rightarrow$ $\nearrow$ $\rightarrow$/$\nearrow$ / ↗
BrightWater  Fewshot–Targeted $\rightarrow$/$\nearrow$ / ↗ $\rightarrow$ $\rightarrow$/ $\searrow$ $\rightarrow$
@@
Table 5: Direction of change with increasing parameter count, by dataset and shot type. Up ()($\uparrow$)( ↑ ), down ()($\downarrow$)( ↓ ), flat ()($\rightarrow$)( → ); slanted arrows (,)($\nearrow$,\,$\searrow$)( ↗ , ↘ ) denote weaker trends.
From a practical perspective, accurate prediction of mode shares is crucial for forecasting applications, requiring models with robust calibration and high instant-level accuracy. Models responsive to targeted few-shot prompting are good candidates for supervised fine‑tuning, provided prompting examples are carefully selected to match the dataset and model. Overall, the pairwise analysis shows the balance between instant-level accuracy and distribution-level calibration to aid the selection models that align with the intended application.

### v2

@@ 3.1.4 Learning Style and Learning Potential of LLMs for Predictive Tasks
Figure 6 shows for each model and dataset the percentage change in weighted F1 when moving from zero-shot to few-shot random and from few-shot random to few-shot targeted. These relative gains expose each model’s capacity to learn from uncurated versus carefully chosen in-context examples.
@@
In the Swissmetro survey, most models improve by 4–15% when given random examples and by an additional 2–27% under targeted examples. Notably, Gemma 3 4B loses performance (–2.3%) with random prompts, a clear case of negative learning likely caused by unrepresentative support sets demonstrated by the strong rebound when supplied with curated examples. By contrast, the Stealth 1.2 7B and Stealth 1.3 7B models display substantial random-shot gains and the largest targeted-shot uplifts (17.2%, 27.3%), highlighting their surprising adaptability and learning potential despite modest parameter counts and baseline performance. On the BrightWater dataset, which is characterized by high behavioural heterogeneity, Mistral 7B leads all models with a random-shot gain of 39.0%, while Stealth 1.2 7B peaks under targeted examples at 36.8%. Only the distilled variants of DeepSeek R1 show slight declines (–5.9%, –0.4%) with random sampling as a result of uncurated examples, which might misalign with their fixed reasoning priors. However, once examples are curated, even these distillates recover and show substantial gains of 22.1–25.9%. For the London PMC dataset, where zero-shot baselines are already strong, Gemma 3 12B exhibits the most dramatic learning curve, with a 56.0% improvement under random examples, but also the largest volatility, dropping -3.4% when switching to targeted. Smaller LLaMA variants, such as Llama 3.2 1B, show negative learning under random (–10.6%) but rebound (+10.6%) with targeted examples, these models lightweight architectures appear more sensitive to example quality but can still benefit from well-aligned support sets.
@@
When averaged across all surveys, the stealth models stand out with the largest sequential gains (18.7, 23%) from zero to random, and an additional 19% to 20.5% from random to targeted, combining strong learning potential with reduced negative learning. On the other hand, the distilled models post modest average improvements (2.4–5.4% from zero to random, 6.6–7.4% from random to targeted), mainly as a result of their limited plasticity once reasoning patterns are fixed during distillation.
@@
These findings support three key insights. First, negative learning can occur when support examples conflict with pretrained expectations or when selected samples are largely misaligned with the test case, particularly in smaller models. Second, instruction-tuned, mid-sized models like those from the Stealth family offer a strong adaptability and learning potential with substantial gains once presented with in-context examples. Third, large, dense transformers such as Gemma 3 12B can leverage random support sets effectively, and learn the underlying relationship between the commuters’ profile and their mode choice, which can be further improved with curated examples. In practice, deploying a few random-shot learning style without quality control should be avoided, especially in complex or noisy environments. Where curated examples are available, their impact can be substantial. Among the tested models, Stealth and Gemma variants emerge as robust and versatile learners, while distilled models remain a reliable choice for zero-shot deployment when examples cannot be supplied.
@@
3.1.5 Prompting Style and Model’s Temperature Performance Impact
Building on the earlier findings regarding model architecture, dataset characteristics, and shot styles, we now turn to a finer grained examination of two additional dimensions influencing predictive performance: the prompt style (Direct vs. Chain-of-Thought/ReAct) and the model temperature (0.5 vs. 1). These two axes offer insight into how LLMs respond to adjustments in prompt structure and sampling stochasticity, independent of parameter count or dataset complexity. These elements, while often overlooked, introduce subtle but meaningful shifts in model behaviour across the three travel survey datasets. Figure 7 shows, for each model across the survey dataset, the percentage point change in weighted F1 score when switching from Chain-of-Thought/ReAct to Direct prompting and from temperature 0.5 to temperature 1, across all three shot styles: zero-shot, few-shot random, and few-shot targeted.
@@ 3.1.5 Prompting Style and Model’s Temperature Performance Impact
Across the board, differences arising from prompt style were larger and more systematic than those from temperature changes. Switching from a Chain-of-Thought/ReAct prompt to a Direct prompt generally improved performance, especially once illustrative examples were available. This effect was most pronounced in the Fewshot_Targetted learning style, where Direct prompting consistently sharpened the model’s focus. For instance, with temperature held at 0.5, using a Direct prompt in the Fewshot_Targetted setting resulted in positive mean F1 score changes across all three datasets (Brightwater: +1.58%, London PMC: +4.63%, Swissmetro: +0.52%) and a positive impact on the weighted F1 score in 9 out of 11 models. However, results in the Zeroshot were noisier and more model dependent; while most models (7) still benefited from concise Direct prompts, others (4) performed better with the reasoning structure provided by CoT/ReAct. These patterns confirm that prompt format is not a one-size-fits-all hyperparameter but must be aligned with the model’s architecture and the specific task context.
@@
Temperature adjustments (from 0.5 to 1.0) produced more subtle effects. Contrary to improving performance, increasing the temperature to 1.0 often led to negligible changes or slight performance declines, especially under a direct prompting style. This finding indicates a general preference for the more deterministic decoding associated with a temperature of 0.5. For example, when using Direct prompting, the number of models that lost performance at $T=1.0$ was greater than the number that gained performance across all datasets and shot types. A critical finding is the non-trivial interaction between temperature and prompt style. The effect of changing temperature often diverged in both magnitude and direction depending on the prompt. For the Qwen 2.5 7B model, increasing temperature under a Direct prompt degraded performance by 3.5-3.5% for few-shot targeted, whereas under a CoT/ReAct prompt, it improved performance by + 5.7%.
@@
Architectural and training differences help explain these heterogeneous responses. Dense instruction-tuned models (e.g., the Gemma, Llama, and Mistral families) typically capitalised on Direct prompts, especially in the Fewshot_Targetted setting, which is more of a reflection of their fine-tuning on concise task specifications. Larger models, like Gemma 3 12B model, although showed consistent improvement with Direct prompting at $T=0.5$, it was overall less sensitive to varying prompting styles or temperatures. On the other hand, smaller models showed extreme sensitivity, with the Llama 3.2 1B model showing a massive +30.37% F1 score increase on the London PMC dataset under these conditions. On the other hand, certain distilled reasoning models, such as DeepSeek R1 Distill Qwen 7B, proved more fragile to prompt compression, often losing accuracy when the CoT/ReAct reasoning chain was removed. These models were also more sensitive to increases in temperature, which often introduced harmful variability.
@@
3.1.6 Comparative Performance on Stated vs. Revealed Preference Datasets
To examine whether LLMs exhibit different learning behaviour when exposed to revealed versus stated preference data, we compare their performance across the Swissmetro and BrightWater stated-preference (SP) datasets and the London PMC revealed-preference (RP) dataset. Across all learning styles, models consistently attain higher weighted F1 and accuracy on the RP dataset, despite its smaller training pool for few-shot examples. Under targeted few-shot prompting, London PMC achieves a mean weighted F1 exceeding 0.67, compared with 0.61 for Swissmetro and 0.58 for BrightWater. Even in the zero-shot setting, the RP dataset produces stronger baselines, indicating that LLMs can infer realistic mode-choice behaviour with minimal contextual guidance. This contrast suggests that causal LLMs generalize better from real-world behavioural signals than from hypothetical survey data. In RP data, observed time–cost trade-offs, accessibility variables, and modal separability provide clear decision boundaries that the model can internalize even from a limited number of examples to choose from. In SP data, however, the presence of hypothetical or aspirational choices introduces additional noise and weakens the relative importance of explanatory features, leading to greater variability and lower per-instance discrimination.
The higher predictive stability of the RP dataset, achieved despite a smaller training pool for few-shot examples, indicates that LLMs can capture genuine travel behaviour more effectively than self-reported intentions. This offers a new perspective on mitigating hypothetical bias in travel survey based modelling, which often limits the realism and practical utility of SP data for planning purposes.
@@ 3.2 Distribution-Level LLMs Mode Choice Predictive Performance Evaluation
Brightwater shows a distinct pattern. Models generally achieve lower instant-level accuracy, yet the distribution-level picture is mixed: DistMAE sits in a tight band (about 0.02–0.12), while JSD and cross-entropy vary widely, with cross-entropy spanning roughly 1.6–11 (Figure 9). Two structural features explain this. First, the choice set contains nine dynamic alternatives, several with small shares. With many thin tails, absolute deviations accumulate slowly: a model can mis-allocate small amounts across rare modes and still keep the sum of absolute errors modest, hence the low DistMAE. JSD responds to relative error: a shift from 0.01 to 0.03 is small in absolute terms but large in relative terms, so JSD rises even when DistMAE stays low. Cross-entropy is stricter still: if a run assigns zero selections to a present class, the loss spikes. Second, the respondent population is comparatively homogeneous. Although the model never observes the aggregate distribution, many instances share similar covariates, so models tend to produce similar selections across cases. When aggregated, the resulting shares sit close to the empirical shares (low DistMAE) but per-instance discrimination remains modest (lower F1).
@@
London PMC achieves the highest instant-level accuracy, with moderate distribution errors and few cross-entropy outliers (Figure 10). Three features contribute. First, the option set contains fewer very low-share classes than Brightwater, so tail-risk omissions are rarer and cross-entropy is steadier. Second, covariates separate the dominant mode clearly, time–cost trade-offs and access variables support a firmer decision boundary, which lifts per-instance accuracy and reduces the need to spread selections thinly. Third, the aggregate shares are stable enough that few-shot exemplars provide useful priors; random few-shot already attains high F1, while targeted few-shot further improves calibration by lifting minority-mode counts and lowering JSD with only small changes in F1. Despite these advantages, the distribution metrics do not collapse to a floor: DistMAE and JSD remain mid-range because high-accuracy models still concentrate selections on the predicted winner and trim secondary modes. Qwen 2.5 7B and Mistral 7B balance this well, sitting near the top in F1 while keeping DistMAE and JSD low. Stealth 1.2 7B offers a complementary profile: slightly lower F1 but the best DistMAE and among the lowest cross-entropy, indicating aggregate shares close to the ground truth. Gemma 3 12B reaches very high F1 under fewshot–random but often places too much weight on the chosen alternative, which in some cases lifts DistMAE and JSDs.
@@
Discrepancies between instant- and distribution-level metrics are expected. A model can score highly on weighted F1 by excelling on majority classes while under‑selecting minority modes. This elevates JSD and, when a present class receives zero selections, produces large cross‑entropy. Conversely, a model that disperses selections more evenly may forgo some top‑1 hits (lower weighted F1) yet produce aggregate shares closer to the truth, which lowers DistMAE. The prompting strategy controls this balance at the instance level. Targeted few‑shot steers selections toward locally plausible alternatives; F1 improves and minority‑mode counts rise, which usually lowers JSD and cross‑entropy. Random few‑shot improves F1 over zero‑shot but can still misalign selections with the test context, leaving JSD elevated and occasionally producing cross‑entropy spikes. Zero‑shot yields both lower F1 and weaker calibration.
@@
In terms of the Model size effects, Figure 11 and Table 3 jointly summarize how parameter count relates to performance. The figure shows points for individual runs and per‑dataset mean trends at each parameter size. The table encodes the dominant direction of change with arrows: up ($\uparrow$), down ($\downarrow$), approximately flat ($\rightarrow$); slanted arrows ($\nearrow,\,\searrow$) indicate weaker tendencies.
@@
BrigthWater   Fewshot–Random $\rightarrow/\nearrow$ $\rightarrow$ $\nearrow$ $\rightarrow/\nearrow$
BrightWater   Fewshot–Targeted $\rightarrow/\nearrow$ $\rightarrow$ $\rightarrow$/ $\searrow$ $\rightarrow$
@@
Table 3: Direction of change with increasing parameter count, by dataset and shot type. Up $(\uparrow)$, down $(\downarrow)$, flat $(\rightarrow)$; slanted arrows $(\nearrow,\,\searrow)$ denote weaker trends.
Second, interaction with learning style. Zeroshot exposes raw inductive bias: when the class structure is skewed (Brightwater), larger models amplify that bias unless guided. Random few‑shot often lifts F1 but does not guarantee coverage of rare modes, so JSD and cross entropy can remain elevated. Targeted few‑shot is the setting where scale reliably converts to better calibration: minority‑mode counts rise, DistMAE and JSD tilt downward, and cross‑entropy spikes disappear.
Third, training recipe versus size. Differences among families matter as much as parameter count on London PMC and, to a lesser extent, on Swissmetro under few‑shot. Qwen 2.5–7B and Mistral 7B achieve high F1 with steady calibration across shots. Reasoning‑distilled models such as DeepSeek R1 Distill models (Llama 8B, Qwen 7B) often have an edge at the same parameter scale on Swissmetro: they separate alternatives cleanly while maintaining coverage of minority modes, yielding high F1 with low DistMAE and JSD and few cross‑entropy outliers. On Brightwater the same decisiveness can overshoot in zeroshot or random few‑shot; targeted prompting restores balance and the models then sit among the best calibrated for their size. On London PMC they remain competitive but not uniformly superior: Qwen 2.5 7B and Mistral 7B deliver comparable accuracy with stable divergence, suggesting instruction tuning shapes performance at least as strongly as parameter count.
From a practical perspective, accurate prediction of mode shares is crucial for forecasting applications, requiring models with robust calibration at the distribution level and high instant-level accuracy. Models responsive to targeted few-shot prompting are good candidates for supervised fine‑tuning, provided prompting examples are carefully selected to match the dataset and model. Overall, the pairwise analysis shows the balance between instant-level accuracy and distribution-level calibration to aid the selection models that align with the intended application.